Hi Guys
Stuck with this one, would appreciate any help:
A dealer purchased a number of equally priced articles, for a total cost of £120.00 She sold all but 1 of them making a profit of £1.50 on each article, with a total revenue of £135.00, How many articles did she purchase? I can get the answer in my head to 11, just need to know how to go about getting there correctly.

Answer by scott8148(6628)

cost per article = 120 / x

revenue = [(120/x) + 1.5] * (x - 1) = 135

120 + 1.5x - (120/x) - 1.5 = 135 ___ 1.5x - (120/x) = 16.5

multiplying by 2x ___ 3x^2 - 240 = 33x ___ x^2 - 11x - 80 = 0

factoring ___ (x + 5)(x - 16) = 0


11 doesn't look so good ...